In your country the corporate tax rate is 20%. Company C has a perpetual, after tax, unlevered cash flow equal to 29,000,000, a return on unlevered equity equal to 9,6% and debt for 120,000,000. The interest rate on the company’s debt is 4,5%. The debt is constant and perpetual. The tax rate is unexpectedly raised to 36%. Assuming that the change in tax rate does not affect the expected returns required by investors, what is the return experienced by shareholders, following the change in the tax rate?

A. 0.00%
B. -16.00%
C. -20.00%
D. -32.00%

The correct answer is C.
Answer #1

Answer:

Correct answer is:

C. - 20.00%

Before change of Tax rate:

After tax, unlevered cash flow equal to = $29,000,000

Corporate tax rate = 20%

Pretax cash flow = $29,000,000 / (1 - 20%) = $36,250,000

Less, Interest on debt ($120,000,000 * 4.5%) = $5,400,000

Cash flow after interest = $30,850,000

Cash flow after tax = $30,850,000 * (1 - 20%) = $24,680,000

After change of Tax rate:

Cash flow after interest = $30,850,000

Cash flow after tax = $30,850,000 * (1 - 36%) = $19,744,000

Change return experienced by shareholders = ($19,744,000 - $24,680,000) / $24,680,000 = - 20.00%

As such option C is correct and other options A, B and D are incorrect.